OpenStudy (anonymous):

The product of 9 and a number is 12 less than three times that number. What is the number?

OpenStudy (anonymous):

NEED HELP PRONTO!

OpenStudy (mathstudent55):

First, choose a variable to represent the number. Let's use x for the number.

OpenStudy (mathstudent55):

If the number is x, then what is "the product of 9 and the number"?

OpenStudy (anonymous):

9 x n = 3 x n - 12. I can make the equation, but I am having trouble solving it.

OpenStudy (mathstudent55):

Ok. Your equation is good. You used n for the variable.

OpenStudy (mathstudent55):

Ok, now let's solve it together.

OpenStudy (mathstudent55):

9n = 3n - 12 We want all n's on the left side, so we subtract 3n from both sides.

OpenStudy (mathstudent55):

9n - 3n = 3n - 3n - 12 That eliminates the 3n from the right side giving us: 6n = -12 Ok so far?

OpenStudy (anonymous):

so you subtract 3n from 9n right?

OpenStudy (anonymous):

sorry, internet crshed.

OpenStudy (anonymous):

I just really need the answer so my dad will stop bugging me.

OpenStudy (mathstudent55):

Yes.

OpenStudy (mathstudent55):

Now that you have 6n = -12 you want n alone. n is being multiplied by 6, so divide both sides by 6 6n/6 = -12/6 n = -2

OpenStudy (anonymous):

omg thank you.

OpenStudy (anonymous):

You don't know how much you've made my day.